# Matched Filtering-Based Channel Estimation for AFDM Systems in Doubly Selective Channels

### v1

@@ VI Conclusions
This paper has studied an MF-based CE scheme for AFDM systems over doubly selective channels, by upholding the key idea of sequentially estimating the parameters of each path. We have derived the I/O relationship of AFDM and analyzed the ICCI, SINR and effective SINR loss for different channels.
We have also studied the path ambiguity problem for FDFD channels that causes severe performance deterioration. To address such a path ambiguity problem, we have introduced a novel pilot pattern across two AFDM transmissions, followed by an MF CE scheme. Specifically, by leveraging the orthogonality and separability between different paths, we have shown that the multipath estimation can be decoupled into low-complexity estimations for individual paths. Consequently, the equivalent shift of each path can be estimated sequentially. In addition, different equivalent shifts corresponding to a path are employed to solve the path ambiguity problem.
To further reduce the search space and redundant computation during fractional parameter estimation, we have proposed the MF-GFS CE scheme that outperforms the proposed MF. We have also extended the proposed MF and MF-GFS CE schemes to the typical broadband systems. The simulation results have shown that the proposed schemes offer more significant advantages over existing CE schemes in terms of improved communication performances and lower complexities.
Appendix A Proof of Theorem 1: I/O Relation of AFDM in DAFT Domain
To analyze the impact of each path, a more detailed I/O relationship is required. Substituting (1), (8) and the DAFT kernel $F_{c_{1},c_{2}}\left(\cdot\right)$ into (10), $y\left(m\right)$ can be expressed as (65). Combining all complex exponential terms yields (66). By grouping the terms involving $\bar{n}$ together, $y\left(m\right)$ can be expressed as
$y(m)=\sum_{n=0}^{N-1}{x\left(n\right)h_{w}\left(m,n\right)}+w\left(n\right),$ (62)
where $h_{w}\left(m,n\right)$ is given by (67) at the top of next page, with
$\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)$ $=\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}})\bar{n}}}$ (64)
$=\frac{1}{N}\frac{e^{-j2\pi(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}})}-1}{e^{-j\frac{2\pi}{N}(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}})}-1}$
Therefore, the proof is complete.
$y\left(m\right)=$ $\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{\sum_{i=1}^{P}{h_{i}\sum_{n=0}^{N-1}{x}\left(n\right)e^{j2\pi(c_{1}\left(\bar{n}-l_{\tau_{i}}\right)^{2}+c_{2}n^{2}+\frac{\left(\bar{n}-l_{\tau_{i}}\right)n}{N})}e^{\frac{j2\pi k_{\nu_{i}}}{N}\left(\bar{n}-l_{\tau_{i}}\right)}}}e^{-j2\pi\left(c_{1}\bar{n}^{2}+c_{2}m^{2}+\frac{\bar{n}m}{N}\right)}+w\left(m\right)$ (65)
$=$ $\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{\sum_{i=1}^{P}{h_{i}\sum_{n=0}^{N-1}{x}\left(n\right)e^{\small{\frac{-j2\pi}{N}}\left(2\bar{n}Nc_{1}l_{\tau_{i}}-Nc_{1}l_{\tau_{i}}^{2}-Nc_{2}n^{2}-n\bar{n}+nl_{\tau_{i}}+\left(-k_{\nu_{i}}\bar{n}+k_{\nu_{i}}l_{\tau_{i}}\right)+\left(Nc_{2}m^{2}+\bar{n}m\right)\right)}}}+w\left(m\right).$ (66)
$h_{w}\left(m,n\right)=$ $\frac{1}{N}\sum_{i=1}^{P}{\sum_{\bar{n}=0}^{N-1}{h_{i}e^{\small{\frac{-j2\pi}{N}}\left(2\bar{n}Nc_{1}l_{\tau_{i}}-Nc_{1}l_{\tau_{i}}^{2}-Nc_{2}n^{2}-n\bar{n}+nl_{\tau_{i}}+\left(-k_{\nu_{i}}\bar{n}+k_{\nu_{i}}l_{\tau_{i}}\right)+\left(Nc_{2}m^{2}+\bar{n}m\right)\right)}}}$ (67)
$=$ $\frac{1}{N}\sum_{i=1}^{P}{h_{i}e^{\small{\frac{-j2\pi}{N}}\left(-Nc_{1}l_{\tau_{i}}^{2}+\left(n+k_{\nu_{i}}\right)l_{\tau_{i}}+Nc_{2}\left(m^{2}-n^{2}\right)\right)}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\left(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right)\bar{n}}}}$
$=$ $\sum_{i=1}^{P}{h_{i}\alpha\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)}.$
Appendix B Proof of Theorem 2: $\boldsymbol{A}_{\mathrm{T}}^{\mathrm{H}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ Can Approximate the Identity Matrix
$\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{1}},k_{\nu_{1}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{2}},k_{\nu_{2}}\right)\right|=$ $\left|\sum_{m=0}^{N-1}{\frac{1}{N}\sum_{n_{1}=0}^{N-1}{e^{\small{\frac{j2\pi}{N}}(m-n_{p}+2Nc_{1}l_{\tau_{1}}-k_{\nu_{1}})n_{1}}}\frac{1}{N}\sum_{n_{2}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}(m-n_{p}+2Nc_{1}l_{\tau_{2}}-k_{\nu_{2}})n_{2}}}}\right|$ (68)
@@ Appendix A Proof of Theorem 1: I/O Relation of AFDM in DAFT Domain
$=$ $\frac{1}{N}\left|\sum_{n_{1}=0}^{N-1}{\sum_{n_{2}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}\left[-n_{p}\left(n_{2}-n_{1}\right)+2Nc_{1}(l_{\tau_{2}}n_{2}-l_{\tau_{1}}n_{1})-\left(k_{\nu_{2}}n_{2}-k_{\nu_{1}}n_{1}\right)\right]}}}}\underset{\delta\left(n_{2}-n_{1}\right)}{\underbrace{\frac{1}{N}\sum_{m=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}m\left(n_{2}-n_{1}\right)}}}}\right|$ (69)
@@
$=$ $\frac{1}{N}\left|\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}\left[2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right]\bar{n}}}}\right|$ (70)
$=$ $\frac{1}{N}\left|\frac{\sin\left\{\pi\left[2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right]\right\}}{\sin\left\{\frac{\pi}{N}\left[2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right]\right\}}\right|$ (71)
$\approx$ $\left|\mathrm{sinc}\left\{2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right\}\right|$ (72)
$\approx$ $\left|\mathrm{sinc}\left\{\left(2\left(\small{\frac{\nu_{\max}}{\varDelta f}}+\xi\right)+1\right)(\tau_{2}-\tau_{1})N\varDelta f-\left(\nu_{2}-\nu_{1}\right)\frac{1}{\varDelta f}\right\}\right|.$ (73)
@@
For convenience, the two columns in $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ are denoted as $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{1}},k_{\nu_{1}}\right)$ and $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{2}},k_{\nu_{2}}\right)$, respectively. $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{1}},k_{\nu_{1}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{2}},k_{\nu_{2}}\right)\right|$ can be expressed as (68)-(73), as shown at the top of this page.
Simply substitute the corresponding terms to obtain (68). To obtain (69), we rearrange the order of the summations in (68) and group together all terms involving $m$. In (69), the summation over $m$ equals $1$ only when $n_{1}=n_{2}=\bar{n}$, and 00 otherwise. Thus, (69) simplifies to (70). Based on the result of the geometric series summation in (70), (71) can be obtained using Euler’s formula. Since $\small{\frac{\pi}{N}}\left(2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right)$ approaches 00, it leads to $\sin\left\{\frac{\pi}{N}\left[2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right]\right\}\approx\frac{\pi}{N}\left[2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\right]$, thus yielding (72). By substituting (2) and (5) into (72), (73) is obtained.
For (72), $\left|\boldsymbol{a}\left(l_{\tau_{1}},k_{\nu_{1}}\right)^{\mathrm{H}}\boldsymbol{a}\left(l_{\tau_{2}},k_{\nu_{2}}\right)\right|=1$ when $l_{\tau_{1}}=l_{\tau_{2}}$ and $k_{\nu_{1}}=k_{\nu_{2}}$. For more general cases, since $2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)\gg 1$ in general, $\left|\boldsymbol{a}\left(l_{\tau_{1}},k_{\nu_{1}}\right)^{\mathrm{H}}\boldsymbol{a}\left(l_{\tau_{2}},k_{\nu_{2}}\right)\right|$ is close to 00. In addition, from (73), an increase in either $\xi$ or $N$ leads to $2Nc_{1}(l_{\tau_{2}}-l_{\tau_{1}})-\left(k_{\nu_{2}}-k_{\nu_{1}}\right)$ becoming greater than $1$, which results in $\left|\boldsymbol{a}\left(l_{\tau_{1}},k_{\nu_{1}}\right)^{\mathrm{H}}\boldsymbol{a}\left(l_{\tau_{2}},k_{\nu_{2}}\right)\right|$ approaching 00 more closely. Therefore, this proof is completed.
Appendix C Proof of Theorem 3: $\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}$ Is a Bounded Unimodal Function
@@
The proof of Theorem 3 can be seen in (74)-(77), as shown at the top of the next page. Since $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)$ and $\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)$ have the same structure, substituting $\boldsymbol{y}_{T}=\sum_{j=1}^{P}{x_{p}h_{i}\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)}$ and considering the orthogonality of $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)$ yields (74). Considering that the $c_{2}$ is small enough, (75) can be obtained. Similar to the derivation process of (68)-(71) in the Appendix B, (75) can be written as (76). Since $\hat{\varUpsilon}_{i}-\varUpsilon_{i}\approx 0$, which leads to (77). In addition, $\gamma-\gamma_{i}\in\left[-1,1\right]$. Therefore, $\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}$ is a bounded unimodal function, with its extremum achieved at $\gamma=\gamma_{i}$. Therefore, this proof is completed.
$\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}$ $\approx\left|x_{p}h_{i}\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\right|$ (74)
$\approx\left|x_{p}h_{i}\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\varUpsilon_{i}+\gamma_{i}\right)\right|$ (75)
$\approx\left|\frac{1}{N}x_{p}h_{i}\frac{\sin\left\{\pi\left[\hat{\varUpsilon}_{i}-\varUpsilon_{i}+\gamma-\gamma_{i}\right]\right\}}{\sin\left\{\frac{\pi}{N}\left[\hat{\varUpsilon}_{i}-\varUpsilon_{i}+\gamma-\gamma_{i}\right]\right\}}\right|$ (76)
$\approx\left|x_{p}h_{i}\right|\sin\mathrm{c}\left(\gamma-\gamma_{i}\right)$ (77)

### v3

@@ VI Conclusions
This paper has studied an MF-based CE scheme for AFDM systems, by upholding the key idea of sequentially estimating the parameters of each path. After analyzing the channel in the DAFT domain, we have proposed two low-complexity methods to construct the channel matrix by exploiting the inherent structural properties of the channel matrix, reducing complexity by at least two orders of magnitude. We have proposed the MF-JE CE scheme, which decouples multipath estimation by exploiting path orthogonality and separability. Furthermore, to further reduce the redundant computation in estimating fractional parameters, the MF-GFS-DE CE scheme was developed. This scheme incorporated the GFS algorithm and decoupled joint DD estimation at the cost of a slight performance loss. Both schemes are extended to the typical broadband systems. The simulation results have shown that the proposed schemes offer more significant advantages over existing CE schemes in terms of improved communication performances and lower complexities.
Appendix A Proof of Theorem 1
To analyze the impact of each path, a more detailed I/O relationship is required. Considering the noise-free version of (11), its corresponding discrete-time signal expression is
$r\left[\bar{n}\right]=\frac{1}{\sqrt{N}}\sum_{i=0}^{P-1}{h_{i}\sum_{n=0}^{N-1}{x\left[n\right]}}e^{\frac{j2\pi}{N}k_{\nu_{i}}\left(\bar{n}-l_{\tau_{i}}\right)}$ (62)
$\times e^{j2\pi\left(c_{2}n^{2}+c_{1}\left(\bar{n}-l_{\tau_{i}}\right)^{2}+\frac{n}{N}\left(\bar{n}-l_{\tau_{i}}\right)+d_{\bar{n},n}\iota_{\tau_{i}}\right)},$
where $d_{\bar{n},n}=\sum_{q=0}^{2Nc_{1}}{\beta_{\bar{n},n}\left(q\right)}$, $\beta_{\bar{n},n}\left(q\right)$ is given by
$\beta_{\bar{n},n}\left(q\right)=\begin{cases}q,&\lfloor\frac{t_{n,q}}{T_{s}}\rfloor+1\leq\bar{n}\leq\lfloor\frac{t_{n,q+1}}{T_{s}}\rfloor\\ 0,&\mathrm{otherwise}\\ \end{cases}.$ (63)
Note that once $N$ and $c_{1}$ are fixed, $d_{\bar{n},n}$ is determined and can thus be pre-calculated.
Substituting (62) and the DAFT kernel $\phi_{m}\left(t\right)$ into (12), The noise-free version of $y\left[m\right]$ can be expressed as (64). Finally, the square bracket in (64) is replaced by $H_{\mathrm{eff}}\left[m,n\right]$, which completes the proof.
@@ Appendix A Proof of Theorem 1
$y\left[m\right]$ $=\sum_{n=0}^{N-1}{x\left[n\right]\left[\sum_{i=1}^{P}{h_{i}\underset{\alpha\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)}{\underbrace{e^{\small{\frac{-j2\pi}{N}}\left(-Nc_{1}l_{\tau_{i}}^{2}+\left(n+k_{\nu_{i}}\right)l_{\tau_{i}}+Nc_{2}\left(m^{2}-n^{2}\right)\right)}}}\underset{\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)}{\underbrace{\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\left(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right)\bar{n}-Nd_{\bar{n},n}\iota_{\tau_{i}}}}}}}\right]}$, (64)
$=\sum_{n=0}^{N-1}{x\left[n\right]H_{\mathrm{eff}}\left[m,n\right]}+w\left[n\right].$ (65)
Appendix B Proof of Theorem 2
$\left|\boldsymbol{a}_{i}^{\mathrm{H}}\boldsymbol{a}_{j}\right|$ $=\left|\sum_{m=0}^{N-1}{\frac{1}{N}\sum_{n_{i}=0}^{N-1}{e^{\small{\frac{j2\pi}{N}}\left((m-n_{p}+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}})n_{i}-Nd_{n_{i},n_{p}}\iota_{\tau_{i}}\right)}\frac{1}{N}\sum_{n_{j}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\left((m-n_{p}+2Nc_{1}l_{\tau_{j}}-k_{\nu_{j}})n_{j}-Nd_{n_{j},n_{p}}\iota_{\tau_{j}}\right)}}}}\right|$ (66)
@@ Appendix B Proof of Theorem 2
$=\frac{1}{N}\left|\sum_{n_{i}=0}^{N-1}{\sum_{n_{j}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}\left[-n_{p}\left(n_{j}-n_{i}\right)+2Nc_{1}(l_{\tau_{j}}n_{j}-l_{\tau_{i}}n_{i})-\left(k_{\nu_{j}}n_{j}-k_{\nu_{i}}n_{i}\right)-N\left(d_{n_{j},n_{p}}\iota_{\tau_{j}}-d_{n_{i},n_{p}}\iota_{\tau_{i}}\right)\right]}}\underset{\delta\left(n_{j}-n_{i}\right)}{\underbrace{\frac{1}{N}\sum_{m=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}m\left(n_{j}-n_{i}\right)}}}}}}\right|.$ (67)
@@
For simplicity, we use $\left|\boldsymbol{a}_{i}^{\mathrm{H}}\boldsymbol{a}_{j}\right|$ to represent $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$. $\left|\boldsymbol{a}_{i}^{\mathrm{H}}\boldsymbol{a}_{j}\right|$ can be expressed as (66)-(67), as shown at the top of this page. Let $\varDelta l_{\tau}=l_{\tau_{j}}-l_{\tau_{i}}$, $\varDelta k_{\nu}=k_{\nu_{j}}-k_{\nu_{i}}$, $\left|\boldsymbol{a}_{i}^{\mathrm{H}}\boldsymbol{a}_{j}\right|$ can be represented as
$\begin{split}\left|\boldsymbol{a}_{i}^{\mathrm{H}}\boldsymbol{a}_{j}\right|&=\frac{1}{N}\left|\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}\left\{\left[2Nc_{1}\varDelta l_{\tau}-\varDelta k_{\nu}\right]\bar{n}-Nd_{\bar{n},n_{p}}\left(\iota_{\tau_{j}}-\iota_{\tau_{i}}\right)\right\}}}}\right|\\ &=\begin{cases}1,&\varDelta l_{\tau}=0,\varDelta k_{\nu}=0\\ \left|\sin\mathrm{c}\left(\varDelta k_{\nu}\right)\right|,&\varDelta l_{\tau}=0,\varDelta k_{\nu}\neq 0\\ \frac{1}{N}\left|\sum_{\bar{n}=0}^{N-1}{\eta_{1}\eta_{2}}\right|,&\varDelta l_{\tau}\neq 0,\varDelta k_{\nu}=0\\ \approx 0,&\varDelta l_{\tau}\neq 0,\varDelta k_{\nu}\neq 0\\ \end{cases},\end{split}$ (68)
where $\eta_{1}=e^{\small{\frac{-j2\pi}{N}2Nc_{1}\varDelta l_{\tau}\bar{n}}}$, $\eta_{2}=e^{\small{j2\pi d_{\bar{n},n_{p}}\left(\iota_{\tau_{j}}-\iota_{\tau_{i}}\right)}}$. Similar to the analysis in (22) and (23), when $\varDelta l_{\tau}$ or $\varDelta k_{\nu}$ is zero, (68) simplifies to a sinc function, peaking at $\varDelta k_{\nu}=0$ or $\varDelta l_{\tau}=0$, respectively. Therefore, when $\varDelta l_{\tau}=\varDelta k_{\nu}=0$ (i.e., $i=j$), the peak of $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$ is $1$. As $\varDelta l_{\tau}$ or $\varDelta k_{\nu}$ increases, $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$ is close to zero. This can also be seen in Fig. 14. Therefore, for any $i\neq j$, $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$ generally satisfies
$\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|\leqslant\epsilon.$ (69)
In general, $\epsilon=\underset{i\neq j}{\max}\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|\ll 1$. Therefore, this proof is completed.
Figure 14: Amplitude of $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$.
Appendix C Proof of Theorem 3
@@
The proof of Theorem 3 can be seen in (70)-(71). substituting $\boldsymbol{y}_{T}=\sum_{j=1}^{P}{x_{p}h_{i}\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)}$ and considering the orthogonality of $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)$, (70) is obtained. From (68), with $\varDelta l_{\tau}=l_{\tau_{i}}-\hat{l}_{\tau_{i}}\approx 0$ and $\bar{k}_{\nu_{i}}\approx\hat{\bar{k}}_{\nu_{i}}$, we obtain (71). In addition, $\kappa_{\nu}-\kappa_{\nu_{i}}\in\left[-1,1\right]$. Therefore, $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\hat{\iota}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$ is a bounded unimodal function, with its extremum achieved at $\kappa_{\nu}=\kappa_{\nu_{i}}$. Therefore, this proof is completed.
$\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\hat{\iota}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$
$\hskip 30.0pt\approx\left|x_{p}h_{i}\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{l}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\right|$ (70)
$\hskip 30.0pt\approx\left|x_{p}h_{i}\right|\mathrm{sinc}\left(\kappa_{\nu}-\kappa_{\nu_{i}}\right).$ (71)